# The unified cross-disciplinary model of the operation of neurons

### v2

@@ 3.1.2 Transient state
Models in neuroscience (as reviewed in [66]) almost entirely ignore these aspects. In our physical model, we see that the measurable membrane potential and current change in the function of the ions’ speed, the concentration, and its time derivative; furthermore, all mentioned quantities depend on the effective potential. As Eq.(3) shows, the "thermodynamic electric field" increases as the temperature increases, and raises the setpoint. As Eq.(27) shows, the current decreases as the temperature increases, in line with the experimental evidence [50, 67]. The amplitude of the action potential is decreased, given that the setpoint has increased and its duration is reduced, given that the current decreases.
@@ 4.1 The physical process
In the resting state, we have a condenser that has some low-conductance resistors (the resting ion channels) and a very low intensity (nearly constant) distributed current flows, under the effect of a slightly varying (nearly constant) membrane voltage. Due to the membrane potential, a current through the always-open channels flows on the dendritic surface and the resting channels’ conductivity represents a drain with sufficient transmission. The resting current practically does not reach the AIS: the ion channels along the current’s path practically "shunt" the current. This is a static state that classic neuroscience assumes: some static current in and some static (leakage) current out; no significant gradient. In this state, the correct model is a condenser with a parallelly switched resistor; an integrator-type $RC$ circuit, see Table 2 (although when approaching the threshold voltage, the AIS plays some role. The smaller gradient changes, such as subthreshold excitations, produce "mini-AP waves" [15], providing direct experimental proof that in this state the parallel $RC$ circuit is not correct). Some perturbations must be counterbalanced, but their amount does not exceed a predefined threshold.
@@
In the transient (perturbed) state, the process variable (the membrane’s potential) exceeds the threshold. That event acts as a fast trigger signal. The amount of charged particles (and so: the membrane’s voltage) suddenly increases, and the created gradient drives a current. When $Na^{+}$ ions rush into the intracellular layer, they roughly increase the overall concentration and the potential in that thin layer. All other ions, including $K^{+}$, also feel a driving force. The targeted membrane potential is set electrically, and the driving gradients may behave controversially during the transient period. The actual voltage gradient may temporarily direct the chemical gradients in opposite directions. Phenomena such as ’over-relaxation’, in this sense, similar to ’hyperpolarization’, may happen.
@@
Our results align with the observation (see caption of 11.22 in [45]), that the significant processes occur in a thin layer of the electrolyte proximal to the membrane surface. The amount of unbalanced ions is in the range of $10^{7}$, and so is the amount of rush-in ions. In addition, those ions on the high-concentration side rush into the low-concentration side and cause a significant change in the membrane potential. Their absolute amount is small compared to the total number of ions in the cell, but it is significant compared to the number of unbalanced ions in that layer. However, the layer itself can also be modeled as having just a few ions under their mutual repulsion on the surface or in a few atomic layers on top of each other, depending on the concentration.
@@
Suppose we assume that at the beginning of an AP a large amount of $Na^{+}$ ions are transferred from the extracellular to the intracellular side. In that case, their concentrations change from $145/15$ to $45/115$ and the corresponding thermodynamic voltage contribution changes from +61[mV]+61\ [mV] to 25[mV]-25\ [mV] (i.e., altogether a 86[mV]86\ [mV] sudden increase in membrane potential). For the $Na^{+}$ ions, the resultant potential changes from +4[mV]+4\ [mV] to 83[mV]-83\ [mV]. That means when the AP begins, the "Na-K pump" stops (if the resulting potential provides the driving force for the exchange pump. The only way for the neuron to remove the excess $Na^{+}$ ions is to generate a current toward the AIS (where the other end of the ion channels remained at the extracellular potential), until the $Na^{+}$-specific driving force disappears. The $K^{+}$-specific driving force, changes from 90[mV]-90\ [mV] to 176[mV]-176\ [mV], i.e., the $K^{+}$ intake gets more intensive, misleading researchers into believing that it causes the observed hyperpolarization. However, as Fig. 3 in [15] demonstrates, it occurs instantly, not with a delay; furthermore, the "leakage current" and the stimulus current are negligible. The low extracellular $K^{+}$ concentration plus the low number of channels in the membrane’s wall do not enable a significant increase in the intracellular $K^{+}$. (Given that complex changes occur, including changes in the electric potential that also change the concentrations, different waves start; the statement is not strictly valid.)
@@
As the current flows, the thermodynamic force decreases, and at some point, the resultant force changes its sign; so does the direction of the current. This is seen as the "condenser current" (and mis-identified as an intense $K^{+}$ current). The ions move slowly, and due to the enormous electrostatic repulsion, their "electrostatic fluid" must be continuous. A special "damped wave" is formed (with all secondary mechanical, optical, etc. effects; see section 4.4) and the condenser current reverses. Electricity interprets the phenomenon as "capacitive current" in the equivalent circuit, and thermodynamics interprets it as changed chemical concentration (or the negative amplitude of the damped elastic oscillation). Our discussion adds that the electric repulsion generates a pressure with a shock wave in the neuron’s closed volume.
@@
Figure 11: A summary of producing neuronal AP as a 3-D figure. The thick solid 3-D line shows how the color-coded voltage gradient drives the membrane potential in the function of the time. The three dashed diagram lines are the projections of the spatial curve to the boundary planes of the cube. These 2-D plots are delivered by the physiology; the figure explains how those parameters are interrelated. The red diagram line shows the voltage gradient produced by the rush-in of the $Na^{+}$ ions, see also the middle inset in Fig. 12. It produces the blue diagram line AP, see also the bottom inset in Fig. 12. The interdependence of the gradient and the voltage it produces leads to Fig. 14, left side, that allows to calculate the thermodynamical characteristics, right side, of neuronal operation.

### v3

@@ 5.1.2 Transient state
Models in neuroscience (as reviewed in [75]) almost entirely ignore these aspects. In our physical model, we see that the measurable membrane potential and current change in the function of the ions’ speed, the concentration, and its time derivative. Furthermore, all mentioned quantities depend on the effective potential. As Eq. (3) shows, the "thermodynamic electrical field" increases as the temperature increases, and raises the setpoint. As Eq.(30) shows that the current decreases as the temperature increases, in line with the experimental evidence [47, 76]. The amplitude of the action potential is decreased, given that the setpoint has increased, and its duration is reduced, given that the current decreases.
The thickness of the charged layer significantly affects the resulting electrical field (i.e., the neuron’s ’setpoint’). Suppose biological fragments are formed and settle on the membrane’s intracellular side due to the increased $\Delta z$. In that case, the resting potential (and the threshold potential) may change, potentially leading to neurological diseases. For example, "in brain tissue from Alzheimer’s disease patients …the most common changes were a proximal shift or a lengthening of the AIS", so, “Structural alterations of the AIS are likely to have an impact on normal neuronal activity” [77]. Lengthening means that the resistance of the AIS increases that changes the parameters of the AP; see Eq.(48). As [78] observed, the setpoint changes with temperature (the "thermodynamic electrical field" changes with the temperature, while the "electrical electrical field" does not, so their crossing point changes), reflected in the dynamics of the AP.
5.2 Controlling neuronal stages
From control theory, it is known that the goal of a controlled system is to govern the application of system inputs to drive the system to a desired state while minimizing delay and overshoot, steady-state error, and ensuring control stability. The neuron implements a controller that monitors the controlled process variable (membrane voltage) and compares it with the reference or setpoint (resting potential). Recall that the geometry and the composition and concentration of electrolytes define the setpoint. The difference between the process variable’s actual and desired values, known as the error signal, represents the actual offset potential. It is applied as feedback to generate a control action that brings the controlled process variable to its setpoint.
5.2.1 Neuron as a PID controller
Biology uses a simple PID controller with proportional, integral, and derivative components. In simple words, it means that the output of the system (the output voltage measurable on the AIS) depends on the input voltage (the present: input current if the resistance is constant), the integrated input current(s) (the past), and their derivatives (the future). Notice that calculating the derivatives of the interdependent parameters needs care, as discussed in [62]. The sum of those summands defines the resulting output voltage. The biological case is more complicated than the technical one because biology works with slow ions, and the components have their temporal dependence (applying the laws created for fast currents needs emulation, as discussed in section 5), furthermore, the summands have several constituents and the neurons handle those current-related constituents autonomously. That is, unlike what is assumed in the classic physiology, the output voltage is not necessarily a simple function of the mentioned input variables; furthermore, it is not sufficient to consider the "present" of the system. The classic models all omit the derivatives’ contributions (a direct consequence of using ’clamping’: in that "frozen" state, the "future" cannot be interpreted) except that of the capacitive current, which is the passive consequence of the input currents. Furthermore, the ’Integrate and Fire’-type models consider the integrated currents only in the ’charge-up’ (’Computing’) period of operation, and they entirely omit that different physical processes are going on in different stages of operation. That is, even the best classic models can not describe neuronal operation completely and correctly.
The gradients are used to adjust the process variable through their positive and negative contributions (corresponding to rising and falling edges), and the different speeds of the thermodynamic and electrical interactions minimize the delay (i.e., provide the maximum operational speed, vital for survival). The steady-state error is minimized by setting the process variable to the reference point using long-term stable parameters (geometry and overall concentration). The low-intensity current through the always-open resting ion channels provides dynamic stability in the steady state.
We considered that the neuron has a stable base state. On the one side, this resting state must be dynamically stabilized for little perturbations using as little energy as possible (and to provide a mechanism when the cell grows, divides, or ages). On the other side, it must be able to restore the state after rough perturbations as quickly as possible, causing short-time transients (when restoring the membrane’s potential after issuing a spike). In both the resting and transient states, the system attempts to return to its balanced state, though the mechanisms required differ.
In its excited state, the system aims to provide an intense output that informs the downstream neurons. This overshoot is initiated by a large number of voltage-gated ion channels (distinct from the non-gated ion channels used to maintain the resting potential) distributed in the neuron’s membrane. The overshoot current flows through those persistently open ion channels, which are concentrated in AIS (which has about two orders of magnitude higher channel density than the wall). The intense slow current produces a condenser-like behavior (capacitive current); the phenomena called "polarization" and "hyperpolarization" (not polarization; instead, a movement of completely separated charge in the surface layer of the electrolyte) of the membrane provide the necessary positive and negative error signals for the controller in the transient state by moving the actual potential value of the membrane above and below the resting potential value. The potentials and the electrical field’s magnitudes depend on the concentration and the geometry (the finite size of the membrane). The ions’ chemical nature comes into play only if the polarizability can differ for different molecules.
The time course of the overall control function $u(t)$ of the general PID controller in function of the error variable $e(t)$ in Fig. 11 is described by
where $T_{i}=\frac{K_{p}}{K_{i}}$ is the time integration constant, $T_{d}=\frac{K_{d}}{K_{p}}$ is the derivative time constant (notice the independent time constants). That is, in the world of a neuron, the "theory of everything" is
$V_{M}^{OUT}(t)=K_{p}\biggl(\underbrace{\overbrace{V_{M}^{IN}(t)}^{proportional}}_{clamping,\ synaptic}+\overbrace{\underbrace{\frac{1}{T_{i}}\int_{0}^{t}V_{M}^{IN}(\tau)d\tau}_{membrane^{\prime}s\ wall}}^{integral\ (parallel\ RC)}+\overbrace{\underbrace{T_{d}\frac{dV_{M}^{IN}(t)}{dt}}}^{derivative\ (serial\ RC)}_{AIS,\ synaptic}\,\biggr)$ (44)
The first term represents the constant effect of the external world (including synaptic inputs, clamping, and excitatory effects on brain tissue). The second term represents the time-averaged contribution, such as charging up the membrane, including clamping current. The voltage is mainly due to current through ion channels in the membrane’s wall; it represents a parallel $RC$ circuit. The third term represents the sum of the gradients due to all effects (the external world, the internal processes, and the outflow through the AIS); it represents a serial $RC$ circuit. The underbraces of the equation identify which components of the neuron contribute the terms of the PID equation. The presence of some terms depends on the neuron’s actual stage; furthermore, the $R$ values in the parallel and serial $RC$ circuit differ (resulting in different time constants).
One can reformulate the HH’s famous Eq. (1) in [14] to the form of the PID equation
$\underbrace{I\times R_{AIS}}_{V_{M}^{OUT}}=\overbrace{\underbrace{I_{i}\times R_{AIS}}_{ionic}}^{proportional}+\overbrace{\frac{1}{\underbrace{C_{M}\times R_{AIS}}_{T_{i}}}\int_{0}^{t}\xcancel{(I_{bio}-I_{clamp})(\tau)}d\tau}^{integral,\ clamping\ cancels\ this}+\overbrace{\underbrace{C_{M}\times R_{M}}_{T_{d}}\frac{dV_{M}}{dt}}^{derivative,\ capacitive}$ (45)
The proportional term comprises the ionic (and external) currents. At the time when Eq. (45) was set up, AIS was not yet known, so HH assumed that in the proportional term, the resistance is identical to the membrane’s resistance implemented by the ion channels in the wall (leading to the ideas of "leakage current" and "resting potential", misleading physiological research.) The experimental conditions canceled the integral term entirely: the negative feedback from clamping precisely counterbalances the neuron’s internally generated current, so the current integrates to zero. (BTW: the "integrate and fire" type models artificially revive the forgotten integral term.) The derivative term assumes that all currents are constants, so the voltage changes only due to the capacitive current. It entirely forgets that gradients are needed to operate an electrical system (leading to the ideas of constant-voltage batteries and magically operated resistors in the equivalent circuits). Although it was known from the beginning that the AP s comprise relatively steep rising and falling edges, it was forgotten that a non-constant current contributes a gradient to the derivative term. HH’s formalism describes only the ’present’ (as can be expected when freezing the state), has no predictive power,nor can it give an account of neuronal memories.
HH considered only the proportional term, plus in the derivative term, the voltage gradient due to the condenser (but not the capacitive current itself, directly leading to the wrong ad-hoc hypothesis of the presence of $K^{+}$ for explaining hyperpolarization). Given that clamping obscures the presence of gradient-like changes in the system, they missed that the input currents also produce a voltage gradient and that the output current through the AIS also produces a gradient. Furthermore, the equation shows that external currents (e.g., the rising and falling edges of step functions) also cause significant changes in the output voltage. Similarly, a "foreign" invasion (changing mechanically the positions of charges by pressure, ultrasound, magnetic pulse, changing the concentration in one of the segments, and so on) generates a change in the position of charges on the neuronal membrane. This way, it generates a voltage gradient (along with pressure and other gradients), which may trigger neuronal spikes. As discussed in section 6.1, subthreshold excitations provide direct experimental evidence that the description of the resting state includes both serial and parallel contributions.
From a control theoretical point of view, the condenser geometry and the ion concentrations set the ’always the same’ value (the set-point; see also Figure 1) and the currents serve to keep or restore that value. After introducing a finite-width membrane into an electrolyte, a potential difference between the two membrane surfaces is created; see Eq. (19). When adjusting the membrane’s potential, we must consider its ground and excited states separately, given that the perturbations in these two cases are vastly different. Nature employs various mechanisms to maintain the ground state and recover it after generating a spike. Although nature’s tools are remarkably similar (channels and pumps; here, for discussing ’downhill’ channels, we must consider the charge-up of different components), the significant difference in current intensity warrants a separate discussion.
When comparing Eq. (46) describing the operation of a simple serial $RC$ oscillator taken from the theory of electricity, Eq. (45) describing HH’s differential equation (having PID in mind), and the complete Eq. (43) describing PID in general, one can see that they describe the same process in different (experimental and mathematical) approximations. As experience shows, Eq. (46) describing the ’net electric’ model provides a sufficiently accurate description of the AP, proving that the dominating effect comes from the derivative term and that the thermodynamic term is linearly proportional to the electrical one. Furthermore, as we emphasized by discussing the physical processes and the mathematical terms, to some measure, the "parallel $RC$ circuit" is also present in the process, although its current amplitude is by two orders of magnitude lower than that of the "serial $RC$ circuit". Furthermore, the time constants of the two circuits are also largely different, making the effect of the parallel circuit unnoticeable. The slow current also plays a role here: the ion currents are ’local’. Given that the channels in the membrane’s wall are always open, the current on its way towards the AIS may flow out through a nearby ion channel, so only a fragment of the ’resting current’ reaches the AIS. Eq. (43) provides a high-accuracy description of the process, but for most practical applications using Eq. (46) provides sufficiently good results, while Eq. (45) is based on an incorrect ’physical model’ and requires a series of incorrect ad-hoc hypotheses to describe observations. Note that the forces have both electrical and thermodynamic components, but they are proportional to each other. That is, although the absolute values of the coefficients in the equations are not accurate, their effects are. Surely, the description of the processes is complicated, but appropriate approximations give sufficiently precise and computationally reasonable results.
5.2.2 Robustness of the control circle
An exciting thought experiment is discussed in [30]. "First, we consider a hypothetical membrane that is permeable only to $Na^{+}$ ions. Suppose that $Na^{+}]_{o}$, the outside or extracellular sodium concentration, is 145mM145\ mM, and $Na^{+}]_{i}$ is 12 mM. Moreover, suppose that initially, there is no membrane potential. The diffusion gradient for $Na^{+}$ favors $Na^{+}$ entry into the cell, and the initial $Na^{+}$ influx carries a charge that builds up on the inside of the cell. This produces a potential (recall that separation of charge produces a potential) across the membrane that impedes further $Na^{+}$ ion movement because the positive charges repel the positively charged $Na^{+}$ ion."
The paper correctly explains that in the case of $Na^{+}$ only, the potential is set up to 66.6mV66.6\ mV. (We quietly add that plus a layer of the corresponding $Cl^{-}$ ions on the other segment, in a similar concentration ratio, builds up; so also, the attraction between positive and negative charges contributes to the effect.) In the absence of our argumentation, it cannot explain why that potential builds up; only why that concentration ratio establishes: the Nernst law alone does not explain the effect. Furthermore, the statement is valid only for a well-defined semipermeable membrane. We can add that the same concentrations develop if initially identical concentrations are present on both sides. The finite width of the membrane creates an electrical potential, a consequence of charge separation, as stated. If there are no ions initially inside, then no charge separation occurs, and therefore, no ion transport begins due to the membrane’s electrical potential. However, the thermal driving force exists, and the started ion transport quickly builds the electrical layer, which effectively helps to establish the resting potential.
This robustness also explains why, during replication (cell division), cells also inherit their operating ability and resting potential; the existing membrane cell continues to grow in the same way. Its voltage does not depend on its size, so it remains constant, and the voltage defined in this way adjusts the concentrations accordingly. Moreover, it explains why, during evolution, cells first formed: a couple of lipids came together to form an imperfect membrane with holes, establishing cellular electricity. As Fig. 9 depicts, the concentrations may adapt to the living conditions (whether living in seawater), but the general operating principles remain the same. The seawater environment enables the use of thinner membranes and lower membrane voltages; however, the same electrical field is used in all cases.
Given that neurons share the external (global) concentrations (defined by the vast amount of ions in the bulk), they provide a firmly fixed offset value to the always-the-same electrical potential. This way, the operations of the individual neurons do not interfere. Furthermore, the internal (local) concentrations can adjust themselves even following a very rough charge perturbation while issuing an AP; moreover, when the living organism is growing.
5.3 Goldman-Hodgkin-Katz potential
It has already been stated, based on experimental evidence, that "the membrane permeability to the ions has nothing to do with the potential generation and the ions’ adsorption on the membrane surface generates the membrane potential"; for a review, see [24]. The idea itself is nonsense: in a balanced state, no ion transport happens, so even without permeability, the balanced state persists; the resting potential has nothing to do with either ions’ mobility or permeability or with ion absorption [28]. Furthermore, there is no idea in GHK about whether the ’setpoint’ (why that specific concentration or potential difference) is present and why the same potential is reset after rough perturbations such as issuing an AP or replicating a cell. The causality is reversed: the potential is a static concept and is set electrically, and the two gradients form the experienced concentrations from the available solvent molecules. For the time course of gradient formation, mobility and permeability play a role, but not in determining concentrations or the resting potential.
In our clear physical picture, the thermodynamic forces on one side and on the other side of the membrane are summed. They counterbalance each other; furthermore, jointly the effect of the neuronal condenser, see Fig. 9. As we emphasized, the concentration gradients are ion-specific. Furthermore, as discussed in connection with Eq.(2), the Nernst equation comprises a per-ion indefinite constant (a potential difference). To calculate a linear combination of terms comprising an arbitrary constant is nonsense, and so is adding absolute concentrations on the different sides of the membrane, or changing the base of the logarithm used in a calculation to match the experimental value. It is not more than number magic. The potential is described by coupled equations as discussed in connection with Eq. (42).
The $Ca^{2+}$ ions do not fit into the GHK picture. One of the reasons why GHK cannot be good is that $Ca^{2+}]$ is omitted. Biologically, it is hard to believe that $Ca^{2+}$ does not participate in the game of life (especially since biology sees the need for $Ca^{2+}$ pumps). Physically, a single concentration on one side of the membrane cannot maintain balance; as the different ion-exchange processes shown in Table 1 and Fig. 9 demonstrate. When adding a new ion to the solution, the sum concentration increases, and so the electrical force increases, forcing the previous elements to find new concentrations on both sides. The appearance of a new chemical element indirectly changes the concentrations of the others (a good example is the role of the negligible amount of $Ca^{2+}$).
@@ 6.1 The physical process
In the resting state, we have a condenser that has some low-conductance resistors (the resting ion channels) and a very low intensity (nearly constant) distributed current flows, under the effect of a slightly varying (nearly constant) membrane voltage. Due to the membrane potential, a current through the always-open channels flows on the dendritic surface and the resting channels’ conductivity represents a drain with sufficient transmission. The resting current practically does not reach the AIS: the ion channels along the current’s path practically "shunt" the current. This is a static state that classic neuroscience assumes: some static current in and some static (leakage) current out; no significant gradient. In this state, an almost correct model is a condenser with a parallelly switched resistor; an integrator-type $RC$ circuit, see Table 2 (although when approaching the threshold voltage, the AIS plays some role. The smaller gradient changes, such as subthreshold excitations, produce "mini-AP waves" [43], providing direct experimental proof that in this state the parallel $RC$ circuit is not correct). Some perturbations must be counterbalanced, but their amount does not exceed a predefined threshold.
@@
In the transient (perturbed) state, the process variable (the membrane’s potential) exceeds the threshold. That event acts as a fast trigger signal. The amount of charged particles (and so: the membrane’s voltage) suddenly increases, and the created gradient drives a current. When $Na^{+}$ ions rush into the intracellular layer, they roughly increase the overall concentration and the potential in that thin layer. All other ions, including $K^{+}$, also feel a driving force. The targeted membrane potential is set electrically, and the driving gradients may behave in unexpected ways during the transient period. The actual voltage gradient may temporarily reverse the direction of the chemical gradients.
@@
Our results align with the observation (see caption of 11.22 in [41]), that the significant processes occur in a thin layer of the electrolyte proximal to the membrane surface. The amount of unbalanced ions is in the range of $10^{7}$, and so is the amount of rush-in ions. In addition, those ions on the high-concentration side rush into the low-concentration side and cause a significant change in the membrane potential (and concentration). Their absolute amount is small compared to the total number of ions in the cell, but it is significant compared to the number of unbalanced ions in that layer. However, the layer itself can also be modeled as having just a few ions under their mutual repulsion on the surface or in a few atomic layers on top of each other, depending on the concentration.
@@
Suppose that at the beginning of an AP a large amount of $Na^{+}$ ions are transferred from the extracellular to the intracellular side. In that case, the $K^{+}/Na^{+}$ concentrations change from $145/15$ to $45/115$ and the corresponding thermodynamic voltage contribution changes from +61[mV]+61\ [mV] to 25[mV]-25\ [mV] (i.e., altogether a 86[mV]86\ [mV] sudden increase in membrane potential). For the $Na^{+}$ ions, the resultant potential changes from +4[mV]+4\ [mV] to 83[mV]-83\ [mV]. That means when the AP begins, the "Na-K pump" stops (if the resulting potential provides the driving force for the exchange pump. The only way for the neuron to remove the excess $Na^{+}$ ions is to generate a current toward the AIS (where the other end of the ion channels remained at the extracellular potential), until the $Na^{+}$-specific driving force disappears. The $K^{+}$-specific driving force, changes from 90[mV]-90\ [mV] to 176[mV]-176\ [mV], i.e., the $K^{+}$ intake gets more intensive, misleading researchers into believing that it causes the observed hyperpolarization. However, as Fig. 3 in [43] demonstrates, it occurs instantly, not with a delay; furthermore, the "leakage current" and the stimulus current are negligible. The low extracellular $K^{+}$ concentration plus the low number of channels in the membrane’s wall do not enable a significant increase in the intracellular $K^{+}$. (Given that complex changes occur, including changes in the electrical potential that also change the concentrations, different waves start; the statement is not strictly valid.)
@@
As the current flows, the thermodynamic force decreases, and at some point the resultant force changes sign; so does the current direction. This is seen as the "condenser current" (and mis-identified as an intense $K^{+}$ current). The ions move slowly, and due to the enormous electrostatic repulsion, their "electrostatic fluid" must be continuous. A special "damped wave" is formed (with all secondary mechanical, optical, etc. effects; see section 6.4) and the condenser current reverses. Electricity interprets the phenomenon as "capacitive current" in the equivalent circuit, and thermodynamics interprets it as changed chemical concentration (or the negative amplitude of the damped elastic oscillation). Our discussion adds that electrical repulsion generates a pressure wave, or shock wave, within the neuron’s closed volume.
@@
Figure 12: A summary of producing neuronal AP as a 3-D figure. The thick, solid 3-D line shows how the color-coded voltage gradient drives the membrane potential over time. The three dashed diagram lines are the projections of the spatial curve to the boundary planes of the cube. The physiology delivers 2-D plots; the figure explains how those parameters are interrelated. The red line shows the voltage gradient produced by the rush-in of the $Na^{+}$ ions, see also the middle inset in Fig. 13. It produces the blue diagram line AP, see also the bottom inset in Fig. 13. The interdependence of the gradient and the voltage it produces leads to Fig. 15, left side, that allows to calculate the thermodynamic characteristics, right side, of neuronal operation.